# Reputation in the Shadow of Exit

### v4

@@ Abstract
I study reputation formation in repeated games where player actions endogenously determine the probability the game permanently ends. Permanent exit can render reputation useless even to a patient long-lived player whose actions are perfectly monitored, in stark contrast to canonical commitment payoff theorems. However, I identify tight conditions for the long-run player to attain their Stackelberg payoff in the unique Markovian equilibria. Along the way, I highlight the role of Markovian strategies in pinning down the value of reputation formation. I apply my results to repeated global games with exit, and obtain new predictions about regime survival and a dynamic foundation for risk dominant selection.
Keywords: Repeated Global Games, Reputation, Regime Change, Endogenous Exit.
@@ 1 Introduction
Economists have long understood the extent to which reputation effects could refine payoffs in complete information repeated games as players become arbitrarily patient. Famously, the seminal work of Fudenberg and Levine (1989) showed a patient long-run player could attain a payoff arbitrarily close to their Stackelberg payoff in the presence of reputational incentives, in stark contrast to folk theorems which could not give a payoff prediction in the absence of those incentives. Of crucial importance to this result is the arbitrary patience assumption: the long-run player expects with high probability to play arbitrarily many periods of the game. This then allows Fudenberg and Levine (1989) to apply their following observation to prove their Stackelberg payoff bound result:
If the long-run player chooses his Stackelberg strategy in every period, there is an upper bound on the number of times the short-run players can fail to play as Stackelberg followers.
This simple observation has proven quite powerful; for very general games which persist with (exogenous) arbitrarily high probability for many periods, much work has been done generalizing this observation to arbitrary monitoring structures and stage game payoffs (see Fudenberg and Levine (1992), Gossner (2011) and Luo and Wolitzky (2025)).
@@
There are, however, many settings where the probability the game itself persists depends on the actions of short-run players. Consider for example the chain-store game of Kreps and Wilson (1982) and Milgrom and Roberts (1982). A crucial assumption in their setup is that the incumbent is able to repeatedly maintain their position, even if they unsuccessfully deter entry. Yet in many markets, successful entry by a challenger can displace an incumbent, changing the structure of the game in future periods or forcing the incumbent themselves to exit. Similarly, in the product choice game of Mailath and Samuelson (2006) (Chapter 15.1) and Pei (2020), it may be that a long enough string of low quality forces the firm to exit. Finally, in repeated models of regime change (see Huang (2017)), a successful currency attack could end the current regime (or lead to the toppling of the central bank). Players, realizing this, may change their behavior, which could in turn affect reputational dynamics.
@@
In such games, the expected number of periods play persists may be significantly shorter than the “upper bound on the number of times the short-run players can fail to play as Stackelberg followers,” at which point it is not clear whether it is still possible for the long-run player to leverage their reputational incentives to secure a high payoff. Indeed, absent strong assumptions on the exit technology itself, it is impossible for economists to make any predictions in general about payoffs, restoring the indeterminancy of the folk theorem even with reputational incentives (see Proposition 4). Despite this, the literature has until now ignored the effect exit may have on the long-run player’s incentives to secure reputation; it has been universally assumed that the game will be exogenously infinitely repeated, so that the logic of the Fudenberg and Levine (1989) upper bound can be applied.
This paper studies reputation formation where a long-run player (choosing $c$) interacts with many short-run players (choosing $a$) who can be either a rational type or a committed type playing the same action. The key modeling innovation is that there is an exit function $f(a,c)$ in each period that determines whether the game continues into the next stage, or whether the game permanently ends. In several important cases studied by the literature—the chain store game, the product choice (trust) game, regime change games—the specific exit assumptions can distort canonical payoffs predictions in a few ways. First, there is no meaningful theory of normalized discounted theory of payoffs with endogenous exit. Lemma 1 shows that so long as the exit probability is nontrivial (i.e. never $1$), then any normalized discount stream converges to 0 in the patient limit. Not to be deterred, this paper resolves this problem by dealing directly with the unnormalized payoff streams. This lends itself to new challenges as well—in particular, Proposition 1 shows that the optimal set of commitment postures varies with the discount rate (albeit, in a monotone way). This makes it unclear whether there can exist a “single” commitment posture that remains fixed as $\delta\to 1$, as is often done to establish payoff predictions in the reputation literature. Instead, it necessitates restrictions on the sets of feasible commitment postures, which are handled in the setup to the main result. Thus, even without studying the payoff set of the repeated game, exit already requires significant modification to the standard analysis.
@@
The main result of this paper is that, in the payoff setting studied by this paper described above, two additional conditions are jointly necessary and sufficient for reputation effects to select the long-run player’s Stackelberg payoff as the unique equilibrium (Markovian) payoff:
@@
    An incentive compatibility condition: the effective discount factor must be such that the long-run player must be willing to take the Stackelberg action when they are maximally rewarded for doing so.
@@
    A credible belief condition: whenever the long-run player has incentives for reputation formation, short-run players must believe that they will pursue them.
@@
The incentive compatibility condition is captured by an inequality that is assumed for the statement of Theorem 1 and ensures the Stackelberg payoff is attainable at some equilibrium in the (complete information) repeated game (see also Proposition 2). The credible belief condition is captured by a Markovian refinement, that short-run player actions are measurable with respect to their belief about the long-run players’ type. This ensures that short-run players act on their beliefs about the long-run player’s rationality, and rules out “revelation equilibria” that have the long-run player reveal rationality today, but which guarantee exit today with high probability. Proposition 4 constructs revelation equilibria of this form which guarantee the long-run player an arbitrarily low payoff absent the Markovian refinement, even when payoffs satisfy my incentive compatibility condition.
@@
Theorem 1 also leads to novel comparisons of Markovian equilibria. In the complete information game, Markovian equilibrium cannot sustain the Stackelberg payoff even when it is an equilibrium of the repeated game in general (Propositions 2 and 3). However, when the regime faces reputational incentives, the unique Markovian equilibrium is the one where the long-run player plays their Stackelberg action at every period and perfectly pools with the committed type [^2] This contrasts with past work, which shows that even under perfect monitoring there can be a continuum of equilibria, see Pei and Li (2021). Thus, the Markovian refinement not only helps select the regime’s best equilibrium payoff, but allows the regime to do strictly outperform, in the presence of reputational incentives, their best Markovian payoff in the complete information game. With perfect monitoring, this phenomena is new, and is a consequence of exit.
@@
Finally, I apply my model to repeated globalized regime change games with exit. A long-run player endogenously determines (at a cost) some coordination friction, after which short-run players see noisy signals about the regime’s strength and decide to attack. The game continues if and only if sufficiently few players attack that the regime is able to overcome the revolution in that period. Such a setup is similar to the game in Angeletos et al. (2007), but with reputational effects and endogenous regime actions. I use Theorem 1 to establish an asymptotic selection result [^3] The asymptotics here require both the idiosyncratic and aggregate noise to vanish, and the regime to become arbitrarily patient in order to satisfy my incentive compatibility condition., Theorem 2: In the unique Markovian equilibrium, the regime attains their Stackelberg payoff if and only if the average strength of the regime is above the endogenously chosen risk-dominant level. This result gives a new dynamic foundation for risk-dominant selection, even when the state is almost known. Conversely, if the average strength is below the risk dominant level, than the regime dies almost surely in the first period and attains an arbitrarily low payoff. Since observed equilibrium play jumps at the risk-dominant threshold, my result sheds light on why seemingly ex-ante similar regimes can differ dramatically ex-post in both (1) observed cruelty and (2) length of survival.
@@
The remainder of this paper is organized as follows. I next review the related literature. Section 2 introduces the model and some basic properties. Section 3 analyzes the Stackelberg action under endogenous exit. Section 4 establishes a folk theorem and a reputation payoff bound in Markov strategies, the main results of the paper. Section 5 applies the framework to study repeated global games of regime change. Section 6 concludes. Omitted proofs are collected in the appendix.
@@ 1.1 Related Literature
This paper contributes to the (extensive) literature on when reputational incentives can select payoffs, started by Kreps and Wilson (1982) and Milgrom and Roberts (1982). I adopt the general framework of the seminal work due to Fudenberg and Levine (1989) and Fudenberg and Levine (1992), who establish the generality of reputational effects in a wide variety of infinitely repeated games with long-run and short-run players (see Fudenberg et al. (1990) and Fudenberg and Levine (1994)). There is a large and rich literature making variations to their model; see Part IV of Mailath and Samuelson (2006) or Mertens et al. (2015) for a survey.
@@
This paper is particularly related to Pei (2024) and Pei (2025), who study endogenous records as a key friction to reputation building. Importantly, they show endogenous records implies there are different roles to the exit probability and patience of the players, and that how these two forces interact can lead to different predictions. I undertake a similar decomposition, but endogenize the exit probability. This paper also relates to Pei and Li (2021), who characterize equilibrium play in the Fudenberg and Levine (1989) model without exit; I characterize unique Markovian play with exit.
@@
Relatedly, Abreu and Gul (2000) and the literature on reputational bargaining (see Fanning and Wolitzky (2022) for a review) assume the game ends when one player concedes, a particular form of exit. While the intuition is similar, the models are quite different: for example, bargaining models are zero-sum, assume long-run long-run players, and have very special exit structures. That these models are economically different can be seen in the new technical challenges that arise in my model and the distinctiveness of the qualitative conclusions that attain.
@@
My paper also relates to models of “bad reputation” (Ely and Valimaki (2003) and Ely et al. (2008)) where a shutdown of statistical learning is similar to exit. I contribute to this literature by first directly working with the exit probability as a primitive instead of a derived object due to monitoring, and show that under this interpretation there are cases where reputation can once again be “good” and benefit the long-run player.
@@
Using Markovian refinements to make predictions in reputational repeated games is not new, and I see my refinement as contributing to a long and historical strand of literature that makes this assumption in order to say something meaningful about equilibrium play. For example, Pei (2016) and Lukyanov et al. (2025) use a Markovian refinement to establish “inverse U-shaped” behavior in labor and goods markets, respectively, to reputational incentives. Giannitsarou and Toxvaerd (2003) use Markovian refinements to render the repeated global game (which is a leading application of my framework) tractable and make predictions. Finally, Bhaskar et al. (2013) provide a foundation for Markov perfection in long-run short-run games with bounded memory: they are the only purifiable equilibrium. In games with reputation, Sannikov and Faingold (2011) show that continuous-time games select a unique sequential equilibrium which is Markovian in the belief of the long-run player’s rationality—exactly my criteria. My paper gives an alternative justification for Markovian criteria, as a minimal way of selecting payoffs in the repeated game with endogenous exit.
@@
Finally, the regime change application builds on the rich literature on equilibrium selection in global games, started by Carlsson and van Damme (1993) and Morris and Shin (1998) (see Morris and Shin (2003) for a survey). Our key innovation in this area is to (1) allow the regime to choose, at a cost, the optimal punishment level, and (2) embed the game into a repeated regime change game reminiscent of Angeletos et al. (2007) and Chassang (2010) but with an i.i.d. state. The second change allows me to analyze reputational effects, and unlike Angeletos et al. (2007) and Angeletos and Pavan (2013), establish the existence of a unique equilibrium. Moreover, the dynamic embedding gives a novel dynamic criteria for risk dominant selection, complementary to the speed of learning interpretation in Koh et al. (2025). The first variation relates to a large literature in global games where some “large” play can take an action that modifies the global game and leads to new applied insights about regime change (see Angeletos et al. (2006), Edmond (2013), and Morris and Shadmehr (2024)). Proposition 1 makes an independent contribution in this direction, by establishing a novel comparative static that rationalizes why stronger or more patient regimes are more likely to be cruel, even when doing so harms the regime.
@@ 2.1 The Stage Game
There are $N+1$ players, $i\in[N]\cup\{L\}$, representing a long-run player $L$ who interacts in each period with $N$ short-run players. Each short-run player $i$ chooses an action from a compact set $a_{i}\in A_{i}$; let $a=\{a_{i}\}\in\prod_{i}A_{i}=A$ denote a profile of short-run player strategies. The long-run player chooses an action $c\in C\subset\mathbb{R}$, where $C$ is a countable (potentially finite) set including 0444In the applications, the set of actions will be a compact interval. Note this is essentially without loss of generality given Assumption 1: If $f:Q\to\mathbb{R}$ is continuous, where $Q\subset X$ is countable and dense, there is a unique continuous function $\bar{f}:X\to\mathbb{R}$ such that $\bar{f}|_{Q}=f$. However, the countability assumption helps significantly when defining equilibrium beliefs off-path, so I maintain it throughout the general analysis.. Short-run player actions generate finite signal $Y$, drawn independently according to a distribution $\rho(\cdot|a)$ [^5] Short-run player monitoring is irrelevant to the analysis and included only for generality.. The long-run player’s action is perfectly monitored.
The action profile $(a,c)$ jointly determines a survival probability, denoted by a function $f:A\times C\to[0,1]$. This can be interpreted as the survival probability of a regime given a short-run action profile $a$ and long-run action $c$.
The long-run player’s expected stage payoff is thus given by $f(a,c)v(a,c)$, where $v(a,c)$ is the payoff the long-run player attains conditional on surviving. Short-run player $i$ has payoffs $u_{i}(a,c)$. I suppose all payoffs are bounded and continuous in all arguments, and moreover normalize long-run payoffs to be nonnegative: $v(a,c)\geq 0$ for all $(a,c)\in A\times C$. One can interpret this normalizing as a requirement that the regime always prefers to live to fight another day: exit is weakly dominated.
@@
$\mathcal{N}(c)=\left\{a:a_{i}\in\operatorname*{\arg\!\max}_{\bar{a}_{i}\in A_{i}}u_{i}(\bar{a}_{i},a_{-i},c)\right\}.$
I make the following assumptions on the survival probability $f$.
@@
$f(a,c)$ is continuous with codomain $(0,1)$. Moreover, $f^{*}(c)=\min_{a\in\mathcal{N}(c)}f(a,c)$ is nondecreasing in $c$.
Continuity is a technical assumption that guarantees the existence of a Stackelberg action for each discount rate. That all survival probabilities are interior can be seen as simply stating that there is always some (potentially arbitrarily small) probability that the regime can survive (or die) regardless of what actions are taken [^6] This helps ensure the likelihood ratio is well-defined. Finally, the requirement $f^{*}(c)$ is nondecreasing in $c$ allows one to interpret the long-run player’s choice $c$ as some action which determines a “probability of exit,” ordered so that larger choices of $c$ imply higher probabilities of survival. Throughout, define
$a^{*}(c)=\min_{a\in\mathcal{N}(c)}\left\{f(a,c)v(a,c)\right\}$
as the set of actions that minimize the long-run player’s stage-game payoff, subject to the fact short-run players are best-replying to each other given $c$ is the long-run player’s action. When there is no loss of ambiguity, I also abuse notation and use $a^{*}(c)$ to mean a selection from the correspondence (for example, $f(a^{*}(c),c)$).
Throughout the analysis, it will be useful (expositionally) to impose the assumption a (myopic) long-run player who need not worry about exit has a dominant action, labeled at $c=0$, and moreover that they would prefer to “milk” their reputation at this dominant action whenever possible.
Assumption 2.
For all $a$, $0=\operatorname*{\arg\!\max}_{c\in C}f(a,c)v(a,c)=\operatorname*{\arg\!\max}_{c\in C}v(a,c)$. For any $c>0$, $v(a^{*}(c),0)>v(a^{*}(0),0)$.
The first part of the assumption implies that the set of payoffs attainable in the one-shot repeated game is exactly $\{f(a,0)v(a,0)\}_{a\in\mathcal{N}(0)}$, and simplifies exposition when constructing punishments for deviating from prescribed equilibrium (though, at the cost of substantially more notational overhead, is not necessary). Let $\bar{V}_{0}(\delta)=\max_{a\in\mathcal{N}(0)}\frac{f(a,0)v(a,0)}{1-\delta f(a,0)}$ denote the long-run player’s largest payoff from playing the static optimum repeatedly, and analogously let $\underline{V}_{0}(\delta)$ be her smallest such payoff. The second part of the assumption guarantees the long-run player strictly benefits from the ability to build a reputation when compared to their worst stage-game Nash payoff. $a^{*}(0)$ can thus be interpreted as the “worst-case” action for the principal, i.e. “exit as often as possible.”

### v5

@@ Abstract
I study reputation formation in repeated games where player actions endogenously determine the probability the game permanently ends. Permanent exit can render reputation useless even to a patient long-lived player whose actions are perfectly monitored, in stark contrast to canonical commitment payoff theorems. However, I identify tight conditions for the long-run player to attain their Stackelberg payoff in the unique Markov equilibrium. Along the way, I highlight the role of Markov strategies in pinning down the value of reputation formation. I apply my results to give qualified commitment foundations for the infinite chain-store game. I also analyze repeated global games with exit, and obtain new predictions about regime survival.
Keywords: Repeated Games, Reputation, Commitment Payoff Theorem, Endogenous Exit.
@@ 1 Introduction
Reputational incentives have long been recognized as a powerful way to refine predictions in infinitely repeated games. By endowing the long-run player with the ability to imitate a commitment type who always takes some action $a$, a rational long-run player can build a reputation for playing $a$ and thus secure the payoff they would obtain as-if they were known to be committed to that action. Successful reputation formation relies on two important features of the repeated game: First, that the game itself is repeated sufficiently often, so that the long-run player can undertake the action enough times to build their reputation, and second, that they are extremely patient, so that the “reputation-building” phase of the game contributes little to their average payoff.
@@
Yet there are a variety of settings where the first assumption may be unrealistic. Consider, for example, the chain store game of Kreps and Wilson (1982) and Milgrom and Roberts (1982) (hereafter, simply “chain store game” or “KWMR”). A crucial assumption in their setup is that the incumbent is able to repeatedly maintain their position in an a-priori fixed number of markets, even if they unsuccessfully deter entry. Yet in many incumbent-entrant games, successful entry by a challenger can displace the incumbent, changing the structure of the game in future periods, and either forcing the incumbent to exit or the game to prematurely end [^2] For example, the monopolist loses their market power in the KWMR chain store game if a challenger successfully enters and is accommodated.. Similarly, in the product choice game of Mailath and Samuelson (2006) (Chapter 15.1) and Pei (2020), it may be that a long enough string of low quality forces the firm to exit. Finally, in models of repeated regime change (c.f. Huang (2017)), a successful currency attack could end the current regime, i.e. lead to the toppling of the central bank. Players, realizing this, may change their behavior, which could in turn affect reputational dynamics.
@@
In such games, the assumption the stage game will be infinitely repeated can be unreasonable, as player actions may lead to premature permanent exit. Consequently, the expected number of periods play persists may be much shorter than the “upper bound on the number of times the short-run players can fail to play as Stackelberg followers,” (c.f. Fudenberg and Levine (1989)) blocking both the long-run player’s ability to build a reputation and for the early stages of the game to be unimportant for their average payoff. Indeed, without strong assumptions on the exit technology itself, little can be said about reputation formation in the shadow of exit, restoring the indeterminancy of the folk theorem present in the complete information analogue (see Proposition 3). Despite the potential appropriateness of directly modeling exit in several of the leading examples studied by the reputation literature, little has been said about the effect exit has on incentives for reputation formation.
This paper studies reputation formation in repeated games where player actions endogenously determine if the game permanently ends. Formally, I consider a model where many short-run players (choosing $a$) interact with a long-run player (choosing $c$) who can be either a rational type or a single committed type who plays a Stackelberg action. In addition to payoffs, the action profile $(a,c)$ also jointly determines a survival probability the game permanently ends. The rational type wishes to maximize their expected payoff by potentially imitating the committed type. Crucially, fear of exit can shut down traditional incentives for reputation formation, as the effective discount rate given short-run player actions can be bounded away from $1$ even when the long-run player is very patient. Thus there is no guarantee that investing in the Stackelberg action today will be rewarded with short-run Stackelberg best replies with high frequency in the future, rendering standard arguments from the reputation literature inapplicable.
@@
The main result of this paper identifies two jointly necessary and sufficient conditions that together are enough for reputation effects to select the long-run player’s Stackelberg payoff as the unique Markov equilibrium payoff:
@@
    Incentive compatibility: the “effective” discount fact is high enough that playing the Stackelberg action can be enforced in some equilibrium of the repeated game.
@@
    Credible beliefs (a Markov refinement): whenever short-run players hold identical beliefs about the long-run player’s rationality, they act the same way.
@@
The incentive compatibility condition is captured by a single enforceability constraint, which is taken as hypothesis for the statement of Theorem 1 (see also Proposition 1). The credible belief condition is captured by a Markov refinement that short-run player actions are measurable with respect to their belief about the long-run players’ type. This ensures that short-run players act on their beliefs about the long-run player’s rationality, and rules out equilibria where the long-run player exits with high probability in the first period due to a lack of variation in continuation play in the future. Indeed, I prove in the absence of this condition an anti-commitment payoff theorem (Proposition 3) by constructing revelation equilibria exactly of this form which guarantee the long-run player an arbitrarily low payoff absent the Markov refinement, even when payoffs satisfy my incentive compatibility condition.
@@
Theorem 1 leads to novel comparisons of Markov equilibria. In the complete information game, Markov equilibrium cannot sustain the Stackelberg payoff even when it is an equilibrium of the repeated game in general (Propositions 1 and 2). However, when the regime faces reputational incentives, the unique Markov equilibrium is the one where the long-run player plays their Stackelberg action at every period and perfectly pools with the committed type. [^3] This contrasts with past work, which shows that even under perfect monitoring there can be a continuum of equilibria, see Pei and Li (2021) Thus, the Markov refinement not only helps select the regime’s best equilibrium payoff, but allows the regime to strictly outperform their best Markov payoff in the complete information game when they have incentives for reputation formation.
I illustrate my results by returning to the chain store game. I show that when accommodating the entrant leads to some probability the game permanently ends, one of two things can happen depending on whether the credible beliefs restriction is adopted. First, in Markov equilibrium, there exists a unique equilibrium profile which secures the long-run player their Stackelberg payoff. However, in sequential equilibrium, the long-run player’s payoff may secure only their minimax payoff, exactly because of pathological equilibria where short-run players vary play in ways that are payoff-irrelevant. That the long-run player cannot secure a high payoff in sequential equilibrium contrasts sharply with the results of the KWMR chain store game absent exit. I interpret these results as giving qualified commitment foundations for the chain store game with exit.
@@
Finally, I apply my model to repeated globalized regime change games with exit. A long-run player endogenously determines, at a cost, some coordination friction, after which short-run players see noisy signals about the regime’s strength and decide to attack. The game continues if and only if sufficiently few players attack that the regime is able to overcome the revolution in that period. Such a setup is similar to the game in Angeletos et al. (2007), but with reputational effects and endogenous regime actions. I use Theorem 1 to establish an asymptotic selection result, [^4] I consider the limit where both the idiosyncratic and aggregate noise vanish in order to satisfy my incentive compatibility condition. Theorem 2: In the unique Markov equilibrium, the regime attains their Stackelberg payoff if and only if the average strength of the regime is above the endogenously chosen risk-dominant level. Conversely, if the average strength is below the risk-dominant level, then the regime dies almost surely in the first period and attains an arbitrarily low payoff. Since observed equilibrium play jumps at the risk-dominant threshold, my result sheds light on why seemingly ex-ante similar regimes can differ dramatically ex-post in both (1) observed punishment and (2) length of survival.
@@
The remainder of this paper is organized as follows. Section 1.1 reviews the related literature. Section 2 introduces the model. Section 3 establishes a folk theorem and a reputation payoff bound in Markov strategies, the main results of the paper. Section 4 applies the model to the chain-store game. Section 5 extends the framework to study repeated global games of regime change. Section 6 concludes. Omitted proofs are collected in the appendix. A supplemental appendix discusses the discounted repeated game.
@@ 1.1 Related Literature
This paper contributes to the extensive literature on when reputational incentives can select payoffs, started by Kreps and Wilson (1982) and Milgrom and Roberts (1982). I adopt the canonical framework of Fudenberg and Levine (1989) and Fudenberg and Levine (1992), who establish the generality of reputational effects in a wide class of infinitely repeated games with long-run and short-run players (see Fudenberg et al. (1990) and Fudenberg and Levine (1994)). There is a large and rich literature studying variations to their model; see Part IV of Mailath and Samuelson (2006) or Mertens et al. (2015) for a survey.
@@
This paper is particularly related to Pei (2024) and Pei (2025), who study endogenous records as a key friction to reputation building. Importantly, Pei shows that endogenous records accord differing roles to the exit probability and patience of players, and consequently there is rich asymptotic behavior depending on which force in the discount rate is taken to $1$. In contrast, I endogenize the exit probability instead of records, so in particular allow it to always be bounded away from its patient limit. This leads to a different exercise than Pei’s models, and thus different predictions despite a similar decomposition of the exit probability. This paper also relates to Pei and Li (2021), who characterize equilibrium play in the Fudenberg and Levine (1989) model without exit; I characterize unique Markovian play with exit.
@@
Relatedly, Abreu and Gul (2000) and the literature on reputational bargaining (see Fanning and Wolitzky (2022) for a review) assume the game ends when one player concedes, a particular form of exit. While the way endogenous exit is encoded is related, the models are quite different: for example, bargaining models are zero-sum, assume all players are long-run, and have very special exit structures. That these models are economically different can be seen in the new technical challenges that arise in my model and the distinctiveness of the qualitative conclusions that attain.
@@
My paper also relates to models of “bad reputation” in extensive form reputation games (see Ely and Valimaki (2003), Ely et al. (2008), and Luo and Wolitzky (2025)) where a shutdown of statistical learning is similar to exit. I contribute to this literature by first directly working with the exit probability as a primitive instead of a derived object due to monitoring, and show that under this interpretation there are cases where reputation can once again secure a high payoff for the long-run player.
@@
Using Markov refinements to make predictions in reputational repeated games is not new, and I see my refinement as contributing to a long and historical strand of literature that makes this assumption in order to say something meaningful about equilibrium play. For example, Pei (2016) and Lukyanov et al. (2025) use a Markov refinement to establish “inverse U-shaped” behavior in labor and goods markets, respectively, to reputational incentives, while Kostadinov and Roldán (2025) focus on Markov equilibrium to characterize the dynamics of optimal monetary policy with reputational concerns. Giannitsarou and Toxvaerd (2003) use Markov refinements to render the repeated global game (which is a leading application of my framework) tractable and make predictions. Bhaskar et al. (2013) provide a foundation for Markov perfection in long-run short-run games with bounded memory: they are the only purifiable equilibrium. In games with reputation, Sannikov and Faingold (2011) show that continuous-time games select a unique sequential equilibrium which is Markovian in the belief of the long-run player’s rationality—exactly my criteria. My paper gives an alternative justification for Markov criteria, as a minimal way of selecting payoffs in the repeated game with endogenous exit.
@@
Finally, the regime change application builds on the rich literature on equilibrium selection in global games, started by Carlsson and van Damme (1993) and Morris and Shin (1998) (see Morris and Shin (2003) for a survey). The model’s key innovation in this area is to (1) allow the regime to choose, at a cost, the optimal punishment level, and (2) embed the game into a repeated regime change game reminiscent of Angeletos et al. (2007) and Chassang (2010) but with an i.i.d. state. The first change relates to a large literature in global games where some “large” player can take an action that modifies the global game and leads to new applied insights about regime change (see Angeletos et al. (2006), Edmond (2013), and Morris and Shadmehr (2024)). Proposition B.1 makes an independent contribution in this direction by establishing a novel comparative static that rationalizes why stronger or more patient regimes are more likely to be cruel, even when doing so harms the regime. The second change allows me to analyze reputational effects, and unlike Angeletos et al. (2007) and Angeletos and Pavan (2013), establish the existence of a unique Markov equilibrium.
@@ 2.1 The Stage Game
There are $N+1$ players, $i\in[N]\cup\{L\}$, representing a long-run player $L$ who interacts in each period with $N$ short-run players. Each short-run player $i$ chooses an action from a compact set $a_{i}\in A_{i}$; let $a=\{a_{i}\}\in\prod_{i}A_{i}=A$ denote a profile of short-run player strategies. The long-run player chooses an action $c\in C\subset\mathbb{R}$, where $C$ is a countable (potentially finite) set including 0.555In the global games application, the set of actions will be a compact interval. Note this is essentially without loss of generality given Assumption 3: If $f:Q\to\mathbb{R}$ is continuous, where $Q\subset X$ is countable and dense, there is a unique continuous function $\overline{f}:X\to\mathbb{R}$ such that $\overline{f}|_{Q}=f$. However, the countability assumption helps significantly when defining equilibrium beliefs off-path, so I maintain it throughout the general analysis. Short-run player actions generate finite signal $Y$, drawn independently according to a distribution $\rho(\cdot|a,c)$. [^6] Short-run player monitoring is irrelevant to the analysis and included only for generality. The long-run player’s action is perfectly monitored.
The action profile $(a,c)$ jointly determines a survival probability, represented by a function $f:A\times C\to[0,1]$. This can be interpreted as the survival probability of a regime given a short-run action profile $a$ and long-run action $c$. The complementarity probability 1f(a,c)1-f(a,c) is the exit probability.
The long-run player’s expected stage payoff is thus given by $f(a,c)v(a,c)$, where $v(a,c)$ is the payoff they attain conditional on surviving. Short-run player $i$ has payoffs $u_{i}(a,c)$. Suppose all payoffs are bounded and continuous in all arguments. Moreover, assume that $f(a,c)<1-\xi$ for some $\xi>0$ always. This can be interpreted as saying there is always some (potentially arbitrarily small) probability exit occurs regardless of what actions are taken. [^7] This ensures the undiscounted payoff streams are well-defined, and in particular converge.
@@
$\mathcal{N}(c)=\left\{a:a_{i}\in\operatorname*{\arg\!\max}_{\overline{a}_{i}\in A_{i}}u_{i}(\overline{a}_{i},a_{-i},c)\right\}.$
Moreover, define
$a^{*}(c)=\operatorname*{\arg\!\min}_{a\in\mathcal{N}(c)}\left\{f(a,c)v(a,c)\right\}$
as the set of actions that minimize the long-run player’s stage-game payoff, subject to the fact short-run players are best-replying to each other given $c$ is the long-run player’s action. When there is no ambiguity, I also abuse notation and use $a^{*}(c)$ to mean a selection from the correspondence (for example, $f(a^{*}(c),c)$).
Throughout the analysis, it will be useful (expositionally) to impose the assumption a (myopic) long-run player who need not worry about exit has a dominant action, labeled at $c=0$, and that they benefit from commitment.
@@
For all $a$, $0=\operatorname*{\arg\!\max}_{c\in C}f(a,c)v(a,c)=\operatorname*{\arg\!\max}_{c\in C}v(a,c)$. Moreover,
$0\not\in\operatorname*{\arg\!\max}_{c\in C}\min_{a\in A}\left\{f(a,c)v(a,c)\right\}$
The first part of the assumption implies the set of payoffs attainable in the one-shot game can be represented as $\{f(a,0)v(a,0)\}_{a\in\mathcal{N}(0)}$, i.e. the profiles where the long-run player chooses 0 and short-run players best reply. This simplifies the exposition by simplifying the “worst” punishments from deviating from a prescribed equilibrium (though at the cost of substantially more notational overhead, is not necessary). Towards this end, let $\overline{V}_{0}=\max_{a\in\mathcal{N}(0)}\frac{f(a,0)v(a,0)}{1-f(a,0)}$ denote the long-run player’s largest payoff from playing the static optimum repeatedly, and analogously let $\underline{V}_{0}$ be their smallest such payoff. The second part of the assumption guarantees guarantees the long-run player strictly benefits from the ability to build a reputation when compared to their worst stage-game Nash payoff.